OpenStudy (anonymous):

How would I solve x(x-3) - 3x(x-4) = -4 * x(x-4) for x?

OpenStudy (anonymous):

so multiply through.

OpenStudy (anonymous):

x(x-3) = x^2-3x

OpenStudy (anonymous):

-3x*x = -3x^2 -3x*-4 = +12x try to do the same for the right side.

OpenStudy (anonymous):

and addd 4 to both sides?

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

My mistake.. its being multiblied by -3x. so i should have 12x=12x and x=1?

OpenStudy (anonymous):

on the left side you have x^2-3x-3x^2+12x on the right side you have -4x(x-4) which is -4x^2+16x so you have x^2-3x-3x^2+12x = -4x^2+16x simplify the left side of the equation -2x^2+9x = -4x^2+16x add 4x^2 to both sides 2x^2 +9x = 16x subtract 9x from both sides 2x^2 = 7x divide both sides by x 2x = 7 x = 7/2 or 0.

OpenStudy (anonymous):

Thank you for breaking it down step by step! greatly appreciated

OpenStudy (anonymous):

you're welcome